Joe's boat can go down stream 10 miles in the same time it goes up stream 8 miles. The current of the stream is 2 mph. How fast does Joes boat go in still water?

Answer by checkley79(3341)

D=RT
10=(R+2)T
T=10/(R+2)
8=(R-2)T
T=8/(R-2)
10/(R+2)=8/(R-2) CROSS MULTIPLY.
8(R+2)=10(R-2)
8R+16=10R-20
8R-10R=-20-16
-2R=-36
R=-36/-2
R=18 MPH FOR THE BOAT IN STILL WATER.
PROOF:
10=(18+2)T
10=20T
T=20/10
T=2 HOURS.
8=(18-2)T
8=16T
T=16/8
T=2 HOURS.